find the x- and y- intercepts of the line. 3x-7y = 42
Answer by mananth(16946)

find the x- and y- intercepts of the line. 3x-7y = 42
...
3x-7y = 42
...
Y intercept plug x=0
0-7y=42
-7y=42
/-7
y=-6
...
x intercept plug y =0
3x-7y=42
3x-0=42
3x=42
/3
x=14
...
m.ananth@hotmail.ca